Solve y^1=(2x+2y+3)/(2x+2y-1)

1
Expert's answer
2022-02-27T16:51:41-0500

"y'=\frac{2x+2y+3}{2x+2y-1}\\
y'=\frac{(2x+2y-1)+4}{2x+2y-1}\\
y'=\frac{2x+2y-1}{2x+2y-1}+\frac{4}{2x+2y-1}\\
y'=1+\frac{4}{2x+2y-1}\\
u(x)=2x+2y-1\\
y=\frac{1}{2}(u-2x+1)\\
y'=\frac{1}{2}(u'-2)\\
\frac{1}{2}(u'-2)=1+\frac{4}{u}\\
\frac{1}{2}u'=\frac{2u+4}{u}\\
\frac{udu}{u+2}=4dx\\
\int{\frac{udu}{u+2}}=\int4dx\\
\int\frac{u+2-2}{u+2}du=\int4dx\\
\int(1-\frac{2}{u+2})du=4x+C\\
u-2ln|u+2|=4x+C\\
2x+2y-1-ln(2x+2y+1)^2=4x+C\\
-2x+2y-1-ln(2x+2y+1)^2=C"